OpenStudy (learningallday):

0.63 – (–9.85) A. –10.48 B. –9.22 C. 9.22 D. 10.48

OpenStudy (anonymous):

D

OpenStudy (learningallday):

ok

OpenStudy (learningallday):

i have a nother one

OpenStudy (anonymous):

okay! :P

OpenStudy (learningallday):

Evaluate d – 6.3 when d = 4.26. A. –10.56 B. –2.04 C. 2.04 D. 10.56

OpenStudy (anonymous):

4.26 - 6.3 = -2.04 B

OpenStudy (learningallday):

Evaluate m – n when m = –1.5 and n = –7.7. A. –9.2 B. –6.2 C. 6.2 D. 9.2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Did you master the assessment?? :D

OpenStudy (anonymous):

or if there is more i'd love to help you with it =)

OpenStudy (learningallday):

multiple choice Which expressions are equal to the expression 5.3 – 14.2? Choose all answers that are correct. A. 5.3 + (–14.2) B. –5.3 + 14.2 C. –14.2 + (–5.3) D. –14.2 – (–5.3)

OpenStudy (anonymous):

hmm, D and A

OpenStudy (anonymous):

Yes, D and A

OpenStudy (anonymous):

so did you master it?

OpenStudy (anonymous):

Did you master it? If not that is okay.